OpenStudy (anonymous):

Write an equation for the hyperbola that satisfies the given set of conditions. Vertices (16,0) and (-16,0) conjugate axis of length 16 units.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what are your thoughts so far?

OpenStudy (anonymous):

honestly I don't know where to start

OpenStudy (amistre64):

start with graphing the verts

OpenStudy (amistre64):

also, start with the basic hyperbola equation to fill in

OpenStudy (anonymous):

I don't know what the hyperbola equation is :/

OpenStudy (amistre64):

then you should look in your notes, or course material since they tell you these things before attempting to give you problems to solve.

OpenStudy (amistre64):

you finding it, and typing it out does more for your learning then me just giving it over to you. which is why i need you to find it for us

OpenStudy (anonymous):

y^2/a^2-x^2/b^2=1?

OpenStudy (amistre64):

thats fine, but we will have to adjust it for the information personal preference is i like to keep as and xs lined up; and bs and ys lined since they are related by the graph

OpenStudy (amistre64):

now, plug in one of your vertexes to see if we need to swap x and y signs

OpenStudy (anonymous):

so 0/a^2-16^2/b^2=1?

OpenStudy (amistre64):

good does -16^2/b^2 = 1 have any real solutions?

OpenStudy (anonymous):

b=16?

OpenStudy (amistre64):

not what im asking, im asking for some basic math knowledge here. -16^2 = b^2 .... but any real number that is ^2 is positive. can a positive number be a negative number at the same time?

OpenStudy (anonymous):

I'm really confused right now

OpenStudy (amistre64):

about what?

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

what are a and b?

OpenStudy (amistre64):

we havent determined that yet. we are trying to determine if the setup is valid: \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\] if its not, then we need to use the other setup \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] which one should we use? well, we are testing your fist idea; test the point (16,0) \[\cancel{\frac{0^2}{a^2}}^{=0}-\frac{16^2}{b^2}=1\color{red}{\implies}-\frac{16^2}{b^2}=1\] can b^2 = -16^2 ? can a positive value ever be a negative value at the same time?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

youll have to show me how ....

OpenStudy (anonymous):

in that equation?

OpenStudy (amistre64):

its the equation we need to use, so yeah

OpenStudy (amistre64):

logically, you cant have 4 apples and NOT have 4 apples you cant have 2 birds and NOT have 2 apples you cant NOT have 10 dollars and have 10 dollars but you say we can :)

OpenStudy (amistre64):

2nd line should be all birds :/

OpenStudy (anonymous):

so the equation is invalid?

OpenStudy (amistre64):

the setup is invalid so we use the alternative \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\] when y=0, x = 16 ... 16^2 = a^2, what is a?

OpenStudy (anonymous):

16

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

now w have to deal with what a conjugate axis is

OpenStudy (amistre64):